# Drift2Matrix: Kernel-Induced Self Representation for Concept Drift Adaptation in Co-evolving Time Series

### v2

@@ Appendix F Detail of Datasets and Experimental setup
For stock datasets, the value of interest we aim to predict is the implied volatility of each stock [^14] We chose to validate our model through forecasting volatility as it provides a quantifiable and objective measure to assess the model’s capability to understand and adapt to market changes, rather than through investment decisions or predicting market trends, which could be influenced by subjective interpretations and external market conditions and fall outside the purview of this study.. Given that true volatility remains elusive, we approximated it using an estimator grounded in realized volatility. We employed the conventional volatility estimator Li & Hong (2011), defined as: $\mathcal{V}_{t}=\sqrt{\sum_{t=1}^{n}(r_{t})^{2}}$, where $r_{t}=\ln(c_{t}/c_{t-1})$ and $c_{t}$ represents the closing price at time $t$. For Stock1, we utilized daily data to gauge monthly volatility, while for Stock2, we used 1-hour intra-day data to determine daily volatility. We conducted online forecasting tests on the Stock2 dataset, segmenting the time series with an 11-day sliding window. Given the constrained intra-day data span (7 months) and our volatility forecasting strategy, extending the sliding window would compromise the available data for assessment. At any time point $t$, the observable data encompasses a period quadruple the window size preceding time $t$, e.g., at time point t=110, our model is trained with $\mathbf{S}\left[66:109\right]$ and forecasts $\mathbf{S}\left[110:121\right]$.
@@
Furthermore, we crafted a Synthetic SyD dataset comprising 500 simulated time series, each generated by a combination of following five nonlinear functions. The synthetic dataset allows the controllability of the structures/numbers of concepts and the availability of ground truth. To make a 780-steps long time series, we randomly choose one of the five functions ten times; every time, this function produces 78 sequential values – which are considered a concept. Table 3 summarizes the statistics of the datasets.
@@
In the kernel representation learning process, we used the Gaussian kernel of the form $\bm{\mathcal{K}}(S_{i},S_{j})=exp(-||S_{i}-S_{j}||^{2}/d_{max}^{2})$, where $d_{max}$ is the maximal distance between series. Parameters $\alpha$, $\gamma$ in Eq. 3 and $\beta$ in Eq. 15 are selected over [2,4,6,8,10,20], [0.1,0.4,0.8,1,4,10] and [5,10,20,40,60,100] respectively and set to be $\alpha=4,\gamma=0.8,\beta=60$ for the best performance. Fig. 6 shows the impact of varying $\alpha$ and $\beta$ on the SyD dataset, while the effect of $\gamma$ can be found in Fig. 17. Table 4 summarizes the computing resources used for our experiments.
Our learning mode of kernel-induced representation can be summarized in two ways depending on the specific scenario:
    When applied to a new time series dataset, Drift2Matrix first re-adaptively determines the most suitable segmentation size ${W_{1},\ldots,W_{b}}$ and learns the concept profiles within each segment through kernel-induced representation. This ensures optimal alignment for concept identification.
    When receiving new data points in an online learning setting (with a fixed segmentation size), Drift2Matrix quickly updates the kernel representation matrix for the new segment, enabling efficient adaptation without recalculating the segmentation size.
@@
Appendix G Parameters Setting of Comparing Methods
For all comparison methods, we set the observable historical data and prediction steps to be the same as for Drift2Matrix. For the ARIMA model, we determined the optimal parameter set using AIC; for the INFORMER model, we configured it with 3 encoder layers, 2 decoder layers, and 8 attention heads, with a model dimension of 512. We trained the model for 10 epochs with a learning rate of 0.001, a batch size of 32, and applied a dropout rate of 0.05; for the N-BEATS model, we adopted three stack modes with 1024 hidden layer units while setting the batch size to 10 for more training examples. For the other comparison methods, we performed fine-tuning on parameters to arrive at the optimal settings for each dataset.
Appendix H Additional Experiments
H.1 Handling Noise and Outliers in Representation Matrix
To evaluate Drift2Matrix’s ability to handle noise, we introduced artificial noise into the SyD dataset. We approached the issue of noise and outliers from two perspectives:
@@ H.1 Handling Noise and Outliers in Representation Matrix
    Noise in the Representation Matrix: Drift2Matrix captures relationships among series, where those belonging to the same concept form a highly correlated submatrix, visible as bright blocks in the heatmap. Self-representation learning naturally identifies noise or outliers, which appear as darker areas due to weak or no correlation. An example of this phenomenon is shown in Fig. 7(a), where a missing connected block can be seen in the lower right corner of the matrix. Importantly, the overall structure of the representation matrix $Z$ remains intact despite the presence of noise.
@@
H.2 The Outputs of Concept Identification with Other Datasets and complete results
@@
H.3 Comprehensive Comparison
In this section, we present a detailed comparison of Drift2Matrix and Auto-D2M with thirteen different models. The results are summarized in Table 7 which includes the performance metrics for each model across all datasets used in our experiments. This comparison highlights the strengths and weaknesses of Drift2Matrix relative to other state-of-the-art models, showcasing its superior ability to handle complex time series data and capture concept drift. Among the seventeen models, ten are forecasting models (ARIMA Box (2013), KNNR Chen & Paschalidis (2019), INFORMER Zhou et al. (2021), N-BEATS Oreshkin et al. (2019), CARD Wang et al. (2023), ETSformer Woo et al. (2022), TimesNet Wu et al. (2022), SparseTSF Lin et al. (2024), FITS Xu et al. (2024), Dlinear Zeng et al. (2023)), the other seven are concept-drift models (MSGARCH Ardia et al. (2019), SD-Markov Bazzi et al. (2017), OrBitMap Matsubara & Sakurai (2019), Cogra Miyaguchi & Kajino (2019), FEDformer Zhou et al. (2022), OneNet Wen et al. (2024) and FSNet Pham et al. (2022)).
Furthermore, we conducted additional Type I and Type II error evaluations for concept detection on the SyD, Stock1, and Stock2 datasets. To contextualize these results, we define **True Positive (TP)**, **False Positive (FP)**, **True Negative (TN)**, and **False Negative (FN)** in the context of our concept detection methodology:
@@ H.3 Comprehensive Comparison
These results demonstrate the robustness of Drift2Matrix in accurately detecting concept drift across datasets with different levels of complexity and domain specificity.
Table 7: Models’ forecasting performance, in terms of RMSE
@@
MSGARCH SD-Markov OrbitMap Cogra FEDformer OneNet FSNet Drift2Matrix Auto-D2M
@@
720 1.093 1.027 0.841 0.863 0.871 0.833 0.864 0.840 0.832
While forecasting series task is not our main focus, we provide a comparison of Drift2Matrix with other models. Results for the extended Auto-D2M, are included but not part of the comparison.
H.4 Forecasting results of N-BEATS on Stock1, Stock2, ETTm2
The forecasting results of N-BEATS are depicted in Fig. 10. Compared to our forecasted result, shown in Fig 2(c) and Fig. 3, N-BEATS falls short in capturing complex concept transitions within multiple time series, revealing the limitations of a model geared solely for single time series forecasting. Additionally, Fig. 11 illustrates visualizations of N-BEATS and Drift2Matrix on the ETTh1 dataset, highlighting the strengths of Drift2Matrix in handling co-evolving time series with concept drift.
H.5 Case Study of Kernel Representation on Motion Segmentation sequences
@@
Figure 11: Visualization of N-BEATS and Drift2Matrix on ETTh1.
H.6 Distribution of RMSE Values Across all Datasets
@@
Figure 13: Results on random four sequences (pepople1, cars10, 1R2TCR, 2T3RTCR) of Hopkins155 database. The top row shows images from the four sequences with superimposed tracked points. The second row is the heatmap of representation matrices yielded by Drift2Matrix. The bottom row is a projection of the representation results into a 3D space.
Figure 14: The heatmap of representation matrices (binarized) learned on SyD by various SOTA and Drift2Matrix.
H.7 Complexity Analysis and Execution Time Evaluation
Solving the optimization problem involves iterative updates to $\mathbf{W}$, $\mathbf{V}$, and $\mathbf{Z}$. The update for $\mathbf{W}$ requires computing the smallest $k$ eigenvalues and eigenvectors of a matrix, with a complexity of $\mathcal{O}(kn^{2})$. Updating $\mathbf{Z}$ involves matrix addition, transposition, and element-wise truncation, resulting in a complexity of $\mathcal{O}(n^{2})$. The update for $\mathbf{V}$, which involves matrix inversion $\left((\bm{\mathcal{K}}+\beta\mathbf{I})^{-1}\right)$, dominates the computation with a complexity of $\mathcal{O}(n^{3})$. Over $T$ iterations, the total computational complexity is $\mathcal{O}(Tn^{3})$, with matrix inversion being the primary bottleneck. The space complexity is $\mathcal{O}(n^{2})$ due to the kernel matrix.
@@ H.7 Complexity Analysis and Execution Time Evaluation
Building on kernel-induced representation learning, we use the initial representation matrix to identify concepts efficiently. Concept prototypes and their associated time series are selected from these concepts to construct $\widehat{\mathbf{S}}$. With $\widehat{\mathbf{S}}$ containing $m(m\ll n)$ time series, we define two kernel matrices: $\widehat{\bm{\mathcal{K}}}\in\mathbb{R}^{m\times m}$ for kernel similarities among the $m$ selected time series, and $\bm{\widetilde{\mathcal{K}}}\in\mathbb{R}^{n\times m}$ for similarities between the entire dataset and the selected time series. Using the Nyström method, the full kernel matrix $\bm{\mathcal{K}}$ is approximated as $\bm{\mathcal{K}}^{*}\approx\bm{\widetilde{\mathcal{K}}}\widehat{\bm{\mathcal{K}}}^{-1}\bm{\widetilde{\mathcal{K}}}^{\mathrm{T}}$. This reduces computational costs significantly as only $\bm{\widetilde{\mathcal{K}}}$ needs computation ($\widehat{\bm{\mathcal{K}}}$ is a subset of $\bm{\widetilde{\mathcal{K}}}$). This approach provides an efficient and scalable solution for large-scale co-evolving time series.
@@
H.8 Ablation Study
H.8.1 Ablation Study on the Number of Series
@@
H.8.2 Ablation Study on Regularizations
We conduct ablation experiments to validate the efficacy of Drift2Matrix’s kernel representation learning. We focus particularly on the regularization and kernelization techniques employed in Eq. 3. Our approach is compared against existing self-representation learning/subspace clustering methods Lu et al. (2018a); Bai & Liang (2020); Elhamifar & Vidal (2013); Ji et al. (2014); Liu et al. (2012) - SSC, LSR, LRR, BDR, EDSC, SSQP, and SSCE, under varying values of $\gamma$. The comparative results are illustrated in Fig. 17. This comparison clearly demonstrates that our model achieves superior performance, outperforming the other methods in terms of RMSE across different datasets for a range of $\gamma$ values.
H.8.3 Ablation Study on Kernel-based methods
We also extend ablation experiments to other kernel-based method to validate the representation capability of Drift2Matrix. We present the representation (binarized version) produced by Drift2Matrix and other SOTA methods on the predicted window of SyD in Fig. 14. Drift2Matrix yields a block diagonal matrix with dense within-cluster scatter and sparse between-cluster separation, revealing the underlying cluster structure. SSC, BDR, SSCE, and LSR perform poorly (i.e., unclearly identified the strong and weak correlations) when the subspaces are nonlinear or overlap. LRR gets improved for those weakly-correlated points (in the light area) but still cannot accurately predict the representation for highly-correlated points (in the dark area), making the cluster undistinguished from each other (Note that undistinguished clusters can lead to bad representations). Although kernel-based methods are adept at handling nonlinear data, they are helpless in the case of potentially locally manifold structures, e.g., SC, KKM, and RKKM fail to distinguish the second cluster from the fourth cluster, and KSSC only finds three clusters.
@@ H.8.3 Ablation Study on Kernel-based methods
H.8.4 Ablation Study on Kernel Functions
In this section, we present an ablation study conducted to evaluate the impact of different kernel functions on the performance of the Drift2Matrix model. This study aims to ascertain the effectiveness of various kernels in capturing nonlinear relationships in time series data.
We explore three different kernel functions in our ablation study: Linear, Polynomial, Sigmoid. These kernels are chosen for their distinct properties in mapping data to higher-dimensional spaces. Each kernel function offers different properties and captures various aspects of the data. The linear kernel is straightforward and effective for linear relationships, while the polynomial kernel can model more complex, non-linear interactions. The sigmoid kernel, inspired by neural networks, and the Gaussian kernel, a popular choice for capturing the locality in data, add more flexibility to the model.

### v3

@@ Appendix F Detail of Datasets and Experimental setup
For stock datasets, the value of interest we aim to predict is the implied volatility of each stock [^13] We chose to validate our model through forecasting volatility as it provides a quantifiable and objective measure to assess the model’s capability to understand and adapt to market changes, rather than through investment decisions or predicting market trends, which could be influenced by subjective interpretations and external market conditions and fall outside the purview of this study.. Given that true volatility remains elusive, we approximated it using an estimator grounded in realized volatility. We employed the conventional volatility estimator (Li & Hong, 2011), defined as: $\mathcal{V}_{t}=\sqrt{\sum_{t=1}^{n}(r_{t})^{2}}$, where $r_{t}=\ln(c_{t}/c_{t-1})$ and $c_{t}$ represents the closing price at time $t$. For Stock1, we utilized daily data to gauge monthly volatility, while for Stock2, we used 1-hour intra-day data to determine daily volatility. We conducted online forecasting tests on the Stock2 dataset, segmenting the time series with an 11-day sliding window. Given the constrained intra-day data span (7 months) and our volatility forecasting strategy, extending the sliding window would compromise the available data for assessment. At any time point $t$, the observable data encompasses a period quadruple the window size preceding time $t$, e.g., at time point t=110, our model is trained with $\mathbf{S}\left[66:109\right]$ and forecasts $\mathbf{S}\left[110:121\right]$.
@@
Furthermore, we crafted a Synthetic SyD dataset comprising 500 simulated time series, each generated by a combination of following five nonlinear functions. The synthetic dataset allows the controllability of the structures/numbers of concepts and the availability of ground truth. To make a 780-steps long time series, we randomly choose one of the following five functions ten times; every time, this function produces 78 sequential values – which are considered a concept. Table 3 summarizes the statistics of the datasets.
@@
In the kernel representation learning process, we used the Gaussian kernel of the form $\boldsymbol{\mathcal{K}}(S_{i},S_{j})=exp(-||S_{i}-S_{j}||^{2}/d_{max}^{2})$, where $d_{max}$ is the maximal distance between series. Parameters $\alpha$, $\gamma$ in Eq. 3 and $\beta$ in Eq. 10 are selected over [2,4,6,8,10,20], [0.1,0.4,0.8,1,4,10] and [5,10,20,40,60,100] respectively and set to be $\alpha=4,\gamma=0.8,\beta=60$ for the best performance. Fig. 7 shows the impact of varying $\alpha$ and $\beta$ on the SyD dataset, while the effect of $\gamma$ can be found in Fig. 17. Table 4 summarizes the computing resources used for our experiments.
Our learning mode of kernel-induced representation can be summarized in two ways depending on the specific scenario: (1) When applied to a new time series dataset, CORAL first re-adaptively determines the most suitable segmentation size ${W_{1},\ldots,W_{b}}$ and learns the concept profiles within each segment through kernel-induced representation. This ensures optimal alignment for concept identification. (2) When receiving new data points in an online learning setting (with a fixed segmentation size), CORAL quickly updates the kernel representation matrix for the new segment, enabling efficient adaptation without recalculating the segmentation size.
@@
Parameters Setting of Comparing Methods For all comparison methods, we set the observable historical data and prediction steps to be the same as for CORAL. For the ARIMA model, we determined the optimal parameter set using AIC; for the INFORMER model, we configured it with 3 encoder layers, 2 decoder layers, and 8 attention heads, with a model dimension of 512. We trained the model for 10 epochs with a learning rate of 0.001, a batch size of 32, and applied a dropout rate of 0.05; for the N-BEATS model, we adopted three stack modes with 1024 hidden layer units while setting the batch size to 10 for more training examples. For the other comparison methods, we performed fine-tuning on parameters to arrive at the optimal settings for each dataset.
Appendix G Additional Experiments
G.1 Handling Noise and Outliers in Representation Matrix
To evaluate CORAL’s ability to handle noise, we introduced artificial noise into the SyD dataset. We approached the issue of noise and outliers from two perspectives:
@@ G.1 Handling Noise and Outliers in Representation Matrix
    Noise in the Representation Matrix: CORAL captures relationships among series, where those belonging to the same concept form a highly correlated submatrix, visible as bright blocks in the heatmap. Self-representation learning naturally identifies noise or outliers, which appear as darker areas due to weak or no correlation. An example of this phenomenon is shown in Fig. 8(a), where a missing connected block can be seen in the lower right corner of the matrix. Importantly, the overall structure of the representation matrix $Z$ remains intact despite the presence of noise.
@@
G.2 The Outputs of Concept Identification with Other Datasets and complete results
@@
G.3 Comprehensive Comparison
In this section, we present a detailed comparison of CORAL and Auto-CORAL with thirteen different models. The results are summarized in Table 7 which includes the performance metrics for each model across all datasets used in our experiments. This comparison highlights the strengths and weaknesses of CORAL relative to other state-of-the-art models, showcasing its superior ability to handle complex time series data and capture concept drift. Among the seventeen models, ten are forecasting models (ARIMA (Box, 2013), KNNR (Chen & Paschalidis, 2019), INFORMER (Zhou et al., 2021), N-BEATS (Oreshkin et al., 2019), CARD (Wang et al., 2023), ETSformer (Woo et al., 2022), TimesNet (Wu et al., 2022), SparseTSF (Lin et al., 2024), FITS (Xu et al., 2024f), Dlinear (Zeng et al., 2023)), the other seven are concept-drift models (MSGARCH (Ardia et al., 2019), SD-Markov (Bazzi et al., 2017), OrBitMap (Matsubara & Sakurai, 2019), Cogra (Miyaguchi & Kajino, 2019), FEDformer (Zhou et al., 2022), OneNet (Wen et al., 2024) and FSNet (Pham et al., 2022)).
Furthermore, we conducted additional Type I and Type II error evaluations for concept detection on the SyD, Stock1, and Stock2 datasets. To contextualize these results, we define True Positive (TP), False Positive (FP), True Negative (TN), and False Negative (FN) in the context of our concept detection methodology:
@@ G.3 Comprehensive Comparison
These results demonstrate the robustness of CORAL in accurately detecting concept drift across datasets with different levels of complexity and domain specificity.
Table 7: Models’ forecasting performance, in terms of RMSE. The best results are in red color and the second best in blue color.
@@
MSGARCH SD-Markov OrbitMap Cogra FEDformer OneNet FSNet CORAL Auto-CORAL
@@
720 1.093 1.027 0.841 0.863 0.871 0.833 0.864 0.840 0.839
While forecasting series task is not our main focus, we provide a comparison of CORAL with other models. Results for the extended Auto-CORAL, are also included.
G.4 Forecasting results of N-BEATS on Stock1, Stock2
The forecasting results of N-BEATS are depicted in Fig. 11. Compared to our forecasted result, shown in Fig 2(c) and Fig. 4, N-BEATS falls short in capturing complex concept transitions within multiple time series, revealing the limitations of a model geared solely for single time series forecasting.
G.5 Case Study of Kernel Representation on Motion Segmentation sequences
@@
G.6 Distribution of RMSE Values Across all Datasets
@@
Figure 13: Results on random four sequences (pepople1, cars10, 1R2TCR, 2T3RTCR) of Hopkins155 database. The top row shows images from the four sequences with superimposed tracked points. The second row is the heatmap of representation matrices yielded by CORAL. The bottom row is a projection of the representation results into a 3D space.
Figure 14: The heatmap of representation matrices (binarized) learned on SyD by various SOTA and CORAL.
G.7 Complexity Analysis and Execution Time Evaluation
Solving the optimization problem involves iterative updates to $\mathbf{W}$, $\mathbf{V}$, and $\mathbf{Z}$. The update for $\mathbf{W}$ requires computing the smallest $k$ eigenvalues and eigenvectors of a matrix, with a complexity of $\mathcal{O}(kn^{2})$. Updating $\mathbf{Z}$ involves matrix addition, transposition, and element-wise truncation, resulting in a complexity of $\mathcal{O}(n^{2})$. The update for $\mathbf{V}$, which involves matrix inversion $\left((\boldsymbol{\mathcal{K}}+\beta\mathbf{I})^{-1}\right)$, dominates the computation with a complexity of $\mathcal{O}(n^{3})$. Over $T$ iterations, the total computational complexity is $\mathcal{O}(Tn^{3})$, with matrix inversion being the primary bottleneck. The space complexity is $\mathcal{O}(n^{2})$ due to the kernel matrix.
@@ G.7 Complexity Analysis and Execution Time Evaluation
Building on kernel-induced representation learning, we use the initial representation matrix to identify concepts efficiently. Concept prototypes and their associated time series are selected from these concepts to construct $\widehat{\mathbf{S}}$. With $\widehat{\mathbf{S}}$ containing $m(m\ll n)$ time series, we define two kernel matrices: $\widehat{\boldsymbol{\mathcal{K}}}\in\mathbb{R}^{m\times m}$ for kernel similarities among the $m$ selected time series, and $\boldsymbol{\widetilde{\mathcal{K}}}\in\mathbb{R}^{n\times m}$ for similarities between the entire dataset and the selected time series. Using the Nyström method, the full kernel matrix $\boldsymbol{\mathcal{K}}$ is approximated as $\boldsymbol{\mathcal{K}}^{*}\approx\boldsymbol{\widetilde{\mathcal{K}}}\widehat{\boldsymbol{\mathcal{K}}}^{-1}\boldsymbol{\widetilde{\mathcal{K}}}^{\mathrm{T}}$. This reduces computational costs significantly as only $\boldsymbol{\widetilde{\mathcal{K}}}$ needs computation ($\widehat{\boldsymbol{\mathcal{K}}}$ is a subset of $\boldsymbol{\widetilde{\mathcal{K}}}$). This approach provides an efficient and scalable solution for large-scale co-evolving time series.
@@
G.8 Ablation Study
G.8.1 Ablation Study on the Number of Series
@@
G.8.2 Ablation Study on Regularizations
We conduct ablation experiments to validate the efficacy of CORAL’s kernel representation learning. We focus particularly on the regularization and kernelization techniques employed in Eq. 3. Our approach is compared against existing self-representation learning/subspace clustering methods (Lu et al., 2018a; Bai & Liang, 2020; Elhamifar & Vidal, 2013; Ji et al., 2014; Liu et al., 2012) - SSC, LSR, LRR, BDR, EDSC, SSQP, and SSCE, under varying values of $\gamma$. The comparative results are illustrated in Fig. 17. This comparison clearly demonstrates that our model achieves superior performance, outperforming the other methods in terms of RMSE across different datasets for a range of $\gamma$ values.
G.8.3 Ablation Study on Kernel-based methods
We also extend ablation experiments to other kernel-based methods (Xu et al., 2022b, 2018, 2020) to validate the representation capability of CORAL. We present the representation (binarized version) produced by CORAL and other SOTA methods on the predicted window of SyD in Fig. 14. CORAL yields a block diagonal matrix with dense within-cluster scatter and sparse between-cluster separation, revealing the underlying cluster structure. SSC, BDR, SSCE, and LSR perform poorly (i.e., unclearly identified the strong and weak correlations) when the subspaces are nonlinear or overlap. LRR gets improved for those weakly-correlated points (in the light area) but still cannot accurately predict the representation for highly-correlated points (in the dark area), making the cluster undistinguished from each other (Note that undistinguished clusters can lead to bad representations). Although kernel-based methods are adept at handling nonlinear data, they are helpless in the case of potentially locally manifold structures, e.g., SC, KKM, and RKKM fail to distinguish the second cluster from the fourth cluster, and KSSC only finds three clusters.
@@ G.8.3 Ablation Study on Kernel-based methods
G.8.4 Ablation Study on Kernel Functions
In this section, we present an ablation study conducted to evaluate the impact of different kernel functions on the performance of the CORAL model. This study aims to ascertain the effectiveness of various kernels in capturing nonlinear relationships in time series data.
We explore three different kernel functions in our ablation study: Linear, Polynomial, Sigmoid (Xu et al., 2022a; Chen et al., 2021). These kernels are chosen for their distinct properties in mapping data to higher-dimensional spaces. Each kernel function offers different properties and captures various aspects of the data. The linear kernel is straightforward and effective for linear relationships, while the polynomial kernel can model more complex, non-linear interactions. The sigmoid kernel, inspired by neural networks, and the Gaussian kernel, a popular choice for capturing the locality in data, add more flexibility to the model.